OpenStudy (anonymous):

Can anybody SHOW me how to do this please. What is the solution to 5a-15+9a=3a+29 Step by step? :) Thanks!

OpenStudy (paki):

combine the like terms and solve.... @cheyennestaggs

OpenStudy (neer2890):

5a+9a-15=3a+29 14a-15=3a+29 14a-3a=15+29 11a=44 a=4